please help

In an ideal 3:17 step-up transformer, the primary power is 34 kW and the secondary current is 30 A. The primary voltage is:


A. 6.25 V

B. 200 V

C. 1133.3 V

D. 6422.2 V

E. 51 V
1
Expert's answer
2018-04-16T09:40:12-0400
Since, the ideal transformer doesn’t lose energy, we can wright:
P=P_P=P_S=I_P V_P=I_S V_S,
here, P_P, P_S is the primary and secondary power, respectively; I_P, I_S is the primary and secondary current, respectively; V_P, V_S is the primary and secondary voltage, respectively.
Then, from this formula we can find the secondary voltage:
P=I_S V_S,
V_S=P/I_S =(34000 W)/(30 A)=1133.3 V.
Since, we have 3:17 step-up transformer, we can write:
V_P/V_S =3/17.
From this formula we can find the primary voltage:
V_P=3/17 V_S=3/17∙1133.3 V=200 V.